For the circuit below, determine,
a. The equivalent resistance.
b. The total current in the circuit.
c. The total power.
d. The current and voltage for R5 and R7
.

Take, V = 48 V, R1 = 25 Ω, R5 = 45 Ω, R6 = 40 Ω, and R7 = 38 Ω.2. 10] For the circuit below, determine, a. The equivalent resistance b. The total current in the circuit. c. The total power
Answer #1

a) R5 and R7 are in series

R = R5 + R7

= 45 + 38 = 83 ohm

83 and 40 ohm are in parallel

R = 83*40 / (83 + 40) = 27ohm

27 and 25 ohm are in series

R(equi) = 27 + 25 = 52 ohm

b) Current in the circuit = V/R = 48/52 = 0.923 A

c) total power = VI = 48*0.923 = 44.3 V

d) current in R5 and R7 = 0.923*40 / (40 + 83) = 0.3A

voltage = IR = 0.923*27 = 24.92 V